A Complete Bibliography of Publications in Biometrika for the decade 1910–1919

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.05

Title word cross-reference

−∞ [Stu17]. β_{1} [Dun11]. β_{2} [Dun11]. η [Pea11b, Pea17]. n [Ber18].

Abac [Her10a]. Aborigines [Tho15]. Acute [Mac11]. Addendum [MP13a]. Additional [Rhi10]. Adequately [Pea14c]. Adjusted [Smi18]. Adoption [Pea19a]. Africa [Mac13]. African [Sta14]. Age [May11]. Albinism [AM10, Pea10b, Sta13]. Albums [Eld19]. Allied [RPB14]. Alternative [Pea10d]. Altitude [AH12]. American [Emm13]. Among [Sta13, Tho19b]. Ancestral [Pea11a]. Ancient [Mac13]. Announcement [Ano14a]. Anomalies [Sta13, Sta14]. Anthropoid [RBP13]. Anthropometric [Sau12, Sch11]. Anthropometrical [Ore15]. Anthropometry [Ano13b, Cra11, Whi15]. Any [PP19, Stu17, Iss18b]. Apes [RBP13]. Appearance [Pea12b]. Appendix [Ano15a, Ano15b, SYC^{+}17]. Application [Gre13, Pea16b, Sno12a]. Appreciation [GP11]. Approximation [Sop13, YP16]. Arising [Har14a]. Artistic [Til18].

Ascertain [E.12]. Association [Pea13b, PH13, Til18, Wai15, Whi15]. Attack [EP15b]. Attack- [EP15b]. Attempt [E.12]. Autumn [TBP11].

B [Gal10, Gor19]. B.N.A. [Ano13c, Smi12]. Bacillus [Pea19a, Hir13]. Based [Emm13]. be [Pea11b, Pea14a, Pea15a, Stu13]. Best [Smi16]. Between [Cla16, E.13, Har11, Har13b, Har14b, Stu17, Wai10]. Bi [Sop14a]. Bi-Serial [Sop14a]. Bibliography [Ano10, Ano13a]. Binomial [Pea15a, Sop14b]. Biology [Pea10a]. Biometric [GW10, Sno12b]. Biometrika [Ano15a]. Biometry [Pea10a]. Birth [E.13]. Biserial [Pea17]. Bloodroot [Har10b]. Body [Lee14]. Bone [Ano13c, Smi12]. Book [E.10a, E.10b, E.13]. Born [Ore15]. Breathing [Bel11]. Breeding [Ano15a, Dav10, Gal10, Pea14a, War17]. Bridge [RBP13]. Broad [Pea13c]. Brownlee [E.13]. Brush [Roc11].

C. [Gal10]. Cairo [Ore15]. Cairo-Born [Ore15]. Calculating [McK12]. Calculation [Har13a]. Can [Pea15a]. Canadensis [Har16, Har10b]. Canaries [Gal12a, Her10b]. Canary [Dav10, Gal10]. Cancer [Cla16, May10, Pea12a]. Capacity [Iss14a, Til18]. Case [AM10, Tur13]. Cases [Gre13, Pea12b, TM12]. Categories [Pea10d, Pea13c]. Cercis [Har16]. Certain [Gre13, Her11, Iss16, Pea14c, Pea15a]. Chance [TM12]. Character [PE13]. Characteristics [Har13b, Har14b, Pea16a]. Characters [Har11, Pea19b, RPB14, Sau12]. Charles [Gor19]. Chatham [Tho15]. Check [Pea19a]. Children [Eld14a, Jon10, Mac11, Wai11]. Choice [Smi18]. Class [Ano17, Har13a, Har14a]. Classes [Har14a]. Claw [MP13b, MP13a]. Clypeal [Lat14]. Coefficient [Eve12, Fis15, Her10c, Her11, Iss18b, Pea13d, Pea15b, RS15, RS18, Sno12a, Sop13, Sop14a, SYC^{+}17, YP16]. Coefficients [Eve19, Har13a, Har14a, Her10a, Iss16, Iss18a, PP19, Pea16a, Tho19b]. Coli [Hir13]. College [Tho15]. Combinations [Har13a]. Comments [And14]. Common [RPB14, RWS^{+}19]. Comparison [TBP11]. Components [Pea10e, Pea15a]. Compound [Pea15a]. Conditions [Pea14a]. Condylo [Ano13c, Smi12]. Condylo-Squamosa [Ano13c, Smi12]. Congenital [Sta14]. Congo [BP12]. Constant [Pea13b]. Constants [Ano13e, Pea10e, Smi16, Smi18]. Construction [E.10b]. Contingency [Pea15b, Pea16c, YP16]. Contours [BP11]. Contribution [Har16, Sim14, Sta13]. Cooperative [RWS^{+}19, TBP11]. Corrected [Cla16, E.13]. Correction [P.12, P.14, Pea11b, Stu13]. Corrections [PP19]. Correlation [And14, Ber18, CP14, Cla16, CK15, Eve10, Eve12, Fis15, GB13, Har11, Har13a, Har14a, Her10a, Her10c, Her11, Iss14b, Iss15, Iss16, Jen12, Ore15, Pea10d, Pea11b, Pea12c, Pea13c, Pea13d, Pea14c, Pea14b, Pea16a, PY18, RS15, RS18, Sno12a, Sop13, Sop14a, SYC^{+}17, Stu13, Stu14, Tho19b]. Correlation-Ratio [Iss15]. Correlations [Lee17]. Corrigenda [Ano13b, Ano13c]. Cow [Pea10c]. Crania [BP12, Tho15]. Cranial [BP11]. Criminal [Whi15]. Criminals [Whi15]. Criteria [PT16]. Criterion [Tho19a]. Cruciferæ [Sim14]. Curve [Gre13]. Curves [Doo17, Pea16b, Tho19a].

Danger [Her11]. Darwinism [Pea10a]. Data [Har16, Iss17, Pea12a, Pea16b]. Davenport [Gal10]. Death [E.13, EP15b, May10, Pea19a, E.13]. Death-rate [Pea19a]. Death-Rates [E.13, EP15b, May10, E.13]. Deathrates [Cla16, PT16]. Deaths [May11]. Decayed [Roc11]. Deduced [Pea10e]. Deformity [MP13b, MP13a]. Degeneracy [Pea10f]. Degree [Tho19b]. Describe [Pea16b]. Described [Pea15a]. Determination [Iss14a, Rhi10]. Determined [TM12]. Determining [Her10a, Iss18a, Lee17, Pea10d]. Development [Har10a]. Deviations [Ano15b, Iss18a, Smi18, Stu19, You16]. Diabetes [Cla16, May11]. Diagram [Rhi10]. Dietaries [Pat14, Pea13e, Pea14e]. Difference [CP14, RS15]. Differential [Pea10f, PT16]. Differentiation [Har14a]. Diphtheria [EP15b]. Discovery [Pea19a]. Disease [Pea12b, TM12]. Disorderly [Har14a]. Distance [Stu17]. Distribution [Ano15b, Fis15, GW10, Gri19, Iss18b, May11, Smi18, SYC^{+}17, TM12]. Distributions [Iss16, PP19, Pea10e, Pea11c, Pea14d, Pea15a, Rhi10, Smi16, Sno12a, Tch18, Tch19]. Dr. [Dav10]. Draba [Sim14]. Drawn [Stu17]. Drones [Lat14]. Due [And14, Stu14]. During [Har10a].

Editorial [Ano15b]. Education [Hos14]. Effect [Mac11, Pea10f]. Eggs [RPB14, RWS^{+}19]. Egypt [Hos14]. Egyptian [Ano13c, Smi12]. Egyptians [Ano13b, Cra11]. Eight [Iss18a]. Elimination [And14, Har10a, Stu14]. Employment [Eld14b]. English [P.15]. Error [Ano17, Her10c, Pea11d, Pea13d, Pea15b, Pea17, RS15, Sop13, Sop14a, YP16]. Errors [Ano13e, Gre13, Her10a, Iss16, Pea14c, Rhi10]. Erythrocytes [AH12]. Especial [Ano13c, Smi12]. Essential [Pea14a]. Estimating [Stu17]. Estimation [Wai11]. Evidence [EP15a]. Examination [AM10, Her14a]. Existence [PT16]. Expectation [Mac13, Tch18, Tch19]. Experimental [Pea16b]. Experiments [Ano15a, War17]. Explanation [Stu19]. Exponential [Sop14b]. Expression [Sop14a]. Extension [Pea14b]. Extent [TM12]. External [Iss14a]. Eyes [Mac11].

F [Ano15a, E.10b]. Factor [Her14a]. Fall [Pea19a]. Fallacious [P.12]. Familial [Tur13]. Family [Coc14]. Female [Emm13]. Femora [P.15]. Fertility [Har11, Llo11, Pea10f, Pea11e]. Fevers [Mac11]. Finding [Eve12]. Finger [Wai15]. Finger-Prints [Wai15]. Finite [PP19]. First [Sch11]. Fisher [Ano15b, SYC^{+}17]. Fit [Pea16b, Tho19a]. Flowers [Sim14]. Fluviatilis [RWS^{+}19]. Fonction [Ber18]. Foot [MP13b, MP13a]. Formula [Iss18b]. Formulae [Her11, Iss14a, Iss18a]. Found [Pea13d]. Fourfold [Eve10, Pea13d]. Fowl [Tur13]. Foxgloves [War17]. Francis [Ano11a]. French [Ber18]. Frequencies [McK12]. Frequency [Ano13e, Ano17, Doo17, Fis15, Gre13, Iss16, Iss18b, PP19, Pea11c, Pea14d, Pea15a, Rhi10, Smi16, Tch18, Tch19]. Front [Ano11b]. Fruit [Har10b].

Fruits [Har10a]. Function [Smi18, Ber18]. Functional [Pea11d]. Functions [Eve10, Lee14]. Further [EP15a, Lee17, Pea11a].

G [E.10b]. Gaboon [BP12]. Galloway [Dav10]. Galton [Ano11a]. Gaussian [Lee14]. General [GP11, Jon10, Pea12c, PE13, Pea16c, TBP11, Wai11]. Generalised [Pea19c]. German [And14]. give [Smi18]. Given [Pea10d, Stu17]. Glasgow [Eld14a]. Goodness [Pea16b, Tho19a]. Goring [Gor19]. Grades [Pea14b]. Greenwood [P.14]. Grouping [Stu13]. Groupings [Eve12, Lee17]. Growth [Jen12]. Guidance [Smi18].

Hair [Mac11]. Hand [Til18]. Hardy [E.10b]. Head [Tur13]. Health [PE13]. Height [Eld14a]. Hereditary [PE13]. Heredity [Las12, Pea11a]. Hibiscus [Har11]. Hierarchical [Tho19b]. High [Dun11, Lee17]. Hispania [Mac13]. Histological [AM10]. History [Eld19]. Homotyposis [Har16, RPB14]. Honduras [Pea13a]. House [Pea12b]. Houses [Pea12a, TM12]. Human [Gal12b, GB13]. Hybridisation [Gal12a].

Identity [Pea14d]. II [Har14b, Iss15, SYC^{+}17]. Illustrations [CP14, Har11]. Increase [AH12]. Indefinitely [Fis15]. Independent [Pea11c, Pea14d]. Index [GW10, Pea10e, Pea11d]. Index-Distributions [Pea10e]. Indian [Emm13]. Individually [Pea15a]. Infantile [Eld14b]. Infinite [PP19]. Influence [EP15b, Hos14, Pea12c, Pea13c]. Information [Ano11c, Ano13f, Ano14b, Ano15c, Ano18]. Inheritance [Her10b, Her14a, Llo11, MP13b, MP13a, Pea10c, Pea11e, Pea19b]. Insanity [Her14a]. Intelligence [Bro10, GP11, Jon10, Wai11, Whi15]. Intensity [Sno13]. Inter [Har13a]. Inter-Class [Har13a]. Intermediates [Hir13]. Internal [Pea10b]. Interpolated [Smi18]. Interpolationsrechnung [E.10a]. Intra [Har13a, Har14a]. Intra-Class [Har13a, Har14a]. Intraoccipitalis [Ano13c, Smi12]. Involucral [Har11]. Iris [Gal12b]. Islands [Tho15]. Isolation [EP15b].

John [E.13]. Junior [P.14].

Karl [Pat14]. Known [MP13b, MP13a].

Laboratory [Sch11]. Large [E.12, Fis15, Har13a, McK12]. Larger [Lee14]. Late [Pea12a]. Law [Pea11a, Pea12a, Stu19, Whi14]. Laws [TM12]. Leaves [Gri19]. Legume [Har16]. Lepidium [Sim14]. Lies [Stu17]. Life [E.13, Eld19, Mac13]. Life-History [Eld19]. Like [Pea10e]. Limit [Sop14b]. Limited [Iss18a, PP19]. Linnæus [Sim14]. Lloyd [Pea11e]. Lobster [MP13b, MP13a]. Lobster-Claw [MP13b, MP13a]. Locality [Wai10]. Lusitania [Mac13].

M.D. [Pea12a]. Macdonell [Ker17]. Made [Pea11b, Pea14c, Emm13, Stu13]. Major [P.14]. Malaria [Wai10]. Malformation [Tur13]. Man [EP15a, RBP13, Sno13]. Manifold [Pea16a]. Markings [Lat14]. Material [Pea15c]. Mathematical [Bro10, Pea11d, Pea19c, Tch18, Tch19]. Matings [Ano15a]. Matter [Ano11b]. Mean [Doo17, Iss18a, Pea15b, Stu17]. Meaning [E.13]. Means [GW10]. Measurement [Pea13c]. Measurements [De 13, Iss14a, Ore15, P.15]. Median [Doo17]. Medieval [P.15]. Mellitus [May11]. Memoir [GW10]. Mendelian [Ano17, Las12, Sno12a]. Mendelism [Her10b]. Method [CP14, McK12, Pea10d, Pea14b, RS15]. Mice [Ano15a]. Milk [Pea10c]. Misstatement [P.14]. Mixed [Iss18a]. Mode [Doo17]. Modern [Ano13b, Cra11, Pea19a]. Moment [Iss18a, Iss18b, PP19]. Moment-Coefficients [PP19]. Moments [Ber18, Har13a, PY18, Tch18, Tch19]. Moriori [Tho15]. Morphology [Har10b]. Mortality [Eld14b, E.10b]. Mosquitoes [Wai10]. Mr [P.14]. Multiple [Iss16, Pea10d, Pea12b, Pea14c, Pea16c, Pea16a]. Museum [Tho15].

N [E.10a]. Nasal [RBP13]. Native [Sta14]. Natives [Ore15, Sta13]. Natural [EP15a, Sno13]. Negro [BP12, Iss14a, May14]. Nest [RWS^{+}19, TBP11]. Nochmals [And14]. Normal [Gre13, Iss18b, PY18, Ber18]. Normale [Ber18]. Note [Ano13d, Dun11, Eld14b, Her14b, May11, May14, Pea10b, Pea10c, Pea11e, Pea13a, Pea13b, Pea14a, RS15, Til18, TM12, You16]. Notes [Gal12b, Hos14]. Novel [Pea16a]. Number [AH12, CK15, Har13a, Iss18b, McK12, Wai10]. Numbers [Whi14]. Numerical [CP14, Iss15]. Nyasaland [Sta13].

Objective [Bro10]. Observational [Pea16b]. Observations [Ano13c, Emm13, Smi12, Smi18, Stu17]. Observed [Smi18]. Obstetric [De 13]. Obstetrician [Ano13d]. Occasionally [Pea14c]. Occipital [Ano13c, Smi12]. Occurrence [McK12]. Occurring [Har10a]. One [Pea10d]. Opinion [Jon10]. Opsonic [GW10, Pea10e, Pea11d]. Order [Iss18b, Tho19b]. Orders [PY18]. Ordinates [PP19]. Osteometric [Pea15c]. Other [Pea10d]. Oxford [Sch11].

Paper [Rhi10]. Papers [Ano15b, SYC^{+}17]. Part [Iss15]. Partial [Her10c, Iss14b, Iss15, Pea16c, Pea16a]. Pearson [Pat14]. Peccavimus [Ano19]. Pelvis [De 13, Emm13]. Perfection [Tho19b]. Persistence [Ano13c, Smi12]. Phagocytosis [GW10]. Phthisis [Pea19a]. Physique [Whi15]. Piebald [Coc14, May14, Pea13a]. Pigmentation [Eld12, Gal12b, Mac11, Sau12, Sta13]. Plant [Har13b, Har14b]. Planted [Har13b, Har14b]. Poisson [Sop14b, Stu19, Whi14]. Polychoric [RS18]. Polynomial [Smi18]. Population [E.12, Fis15, Iss18a, Pea11c, Pea14a, Pea14d, Stu17, TBP11]. Position [And14, Stu14]. Possible [Har13a]. Posterior [Ano13c, Smi12]. Power [Mac11]. Practice [Stu19]. Prediction [De 13]. Preliminary [Til18, War17]. Prevalence [E.12]. Prints [Wai15]. Probabilities [McK12]. Probability [Pea11c, Pea14d, Stu17]. Probable [Ano13e, Ano17, Her10a, Her10c, Iss16, Pea13d, Pea15b, Pea17, RS15, Sop13, Sop14a, YP16]. Problem [Har16, Pea15c]. Produced [Har13b, Har14b]. Product [Iss18b, PY18]. Product-Moment [Iss18b]. Product-Moments [PY18]. Products [Iss18a]. Professor [Pat14, Pea11e]. Proof [P.12]. Proper [Smi18]. Properties [Pea16a]. Provinces [Mac13]. Psychical [Pea19b]. Psychophysical [Tho19a]. Pulse [Bel11, Whi15].

Quadrature [Eve19]. Quality [Pea10c]. Quantitative [Har10b]. Quantity [Pea10c]. Queens [Lat14, TBP11].

R [Ano15a, Ker17]. R. [Ano15b, SYC^{+}17]. Race [Mac11, Sta14]. Racial [Las12]. Random [Gre13, Pea14a]. Range [PP19]. Ranks [Pea14b]. Rapidity [Til18]. Rate [Wai10, Pea19a]. Rates [Bel11, E.13, EP15b, May10, E.13]. Ratio [Iss14b, Iss15, Pea10e, Pea11b, Stu13]. Rays [CK15]. Ready [Ano13d]. Really [Pea11c, Pea14d]. Reckoner [Ano13d]. Records [Ano15a]. Recuperative [Mac11]. Reference [Ano13c, BP12, De 13, GW10, Pea10e, Pea14d, Pea16c, Smi12]. Regard [Pea14c]. Regression [Iss16, Pea16b]. Rejoinder [Gal10, Pea14e]. Relation [Bel11, Doo17, Eld12, Sau12, Wai10]. Relationship [E.13, Har13b, Har14b, RBP13]. Relationships [E.13]. Remarks [Pea11a, Pea11e]. Reply [Pat14]. Report [War17]. Representation [Iss17]. Reproductive [Her14b]. Respiration [Whi15]. Results [Sch11]. Review [E.10a, E.10b, E.13]. Reviewers [Sno12b]. Rome [Mac13]. Royal [Tho15].

S. [RWS^{+}19]. Same [Pea11c, Pea12b, Pea14d]. Sample [Stu17]. Samples [Ano15b, Fis15, GW10, Iss18a, Pea11c, Pea14d, SYC^{+}17, You16]. Sampling [Gre13]. Sanguinaria [Har10b]. Sassafras [Gri19]. School [Eld14a, Jon10, Wai11]. Second [GB13, Sop13, GW10, Rhi10]. Seed [Har13b, Har14b]. Segregation [Lee15]. Selection [EP15a, Her14b, Hir13, Mac11, Pea12c, Sau12, Sno13]. Selective [Har10a]. Separate [Pea10c]. Serial [Sop14a]. Series [Ano13c, Pea15a, Smi12]. Sexing [Pea15c]. Sheppard [P.12]. Should [Pea14a]. Significance [GP11]. Simplified [McK12]. since [Pea19a]. Single [TBP11]. Skew [Iss16]. Skull [BP12, Iss14a]. Skulls [Ano13c, Smi12]. Slopes [PP19]. Small [Ano15b, SYC^{+}17, Whi14]. Somatic [Har11]. Some [CK15, Her14a, Pea10a, Pea16a, War17]. Space [Stu14, And14]. Special [BP12, De 13, GW10, Pea10e, Pea14d, Pea16c]. Specimens [Emm13]. Split [MP13b, MP13a]. Split-Foot [MP13b, MP13a]. Spurious [And14, Har14a, Stu14]. Squamosa [Ano13c, Smi12]. Square [Pea15b]. Squaw [Emm13]. Stable [Pea14a]. Standard [Ano15b, Smi18, You16].

Staphylea [Har10a]. State [Pea14a]. Statistical [Hos14, Iss17, May10, Pat14, Pea13e, Pea14e, Sim14, Sno12b]. Statistics [Pea19c]. Stature [Bel11, Eld12]. Steadiness [Til18]. Strains [Pea14d]. Student [Ano15b, SYC^{+}17]. Studied [Pea14c]. Studies [Her14a, E.13]. Study [BP12, Bro10, Emm13, GW10, GB13, Har10b, Her10b, May10, Pat14, Pea13e, Pea14e, RWS^{+}19, RBP13, Sim14, Sta13, TBP11, Tho15, Wai10, Whi15]. Subject [Pea14c]. Substitutes [Her11]. Suffering [Mac11]. Suggested [Her11]. Suitable [Gre13]. Supplementary [Eve12, Lee17]. Surface [Pea13b, PY18]. Surgeons [Tho15]. Susceptibility [Mac11]. Synchondrosis [Ano13c, Smi12]. Syphilis [E.12].

T [E.10a]. Table [E.13, Lee14, Pea13d, RS18]. Tables [E.10b, Eve10, Eve12, Lee17, Pea16b, Rhi10, Sop14b, Stu17]. Tail [Lee14]. taken [Iss18a]. Tchebycheff [Pea19c]. Teacher [Jon10, GP11, Wai11]. Teeth [Roc11]. Temperature [Whi15]. Terminals [PP19]. Tern [RPB14, RWS^{+}19]. Test [Pea16b]. Tetrachoric [Eve10, Eve12, Lee17]. Th. [Pea12a]. their [Bel11]. Theorems [Pea19c]. Theoretical [Pea16b]. Theories [PH13]. Theory [E.10b, Pea12c, Pea16c, Pea19c]. there [PP19]. Thiele [E.10a]. Third [TBP11]. Those [Pea14c]. Three [You16]. Time [And14, Stu14]. Tooth [Roc11]. Tooth-Brush [Roc11]. Treatment [Pea19a]. Trout [Jen12]. Trypanosome [Pea14d]. Tubercle [Pea19a]. Tuberculosis [Lee15]. Two [Iss18a, Pea11c, Pea14d, PY18, You16]. Type [BP11, Rhi10]. Type-Contours [BP11]. Types [Pea15a].

Umbel [CK15]. Umbelliferae [CK15]. Unequal [McK12]. Unique [Stu17]. Universe [Pea16a]. Urban [E.12]. Used [Pea16b].

Value [Jon10]. Values [Dun11, Fis15, Har14a, Iss18a, Smi16, Smi18]. Variability [GB13, Jen12]. Variable [Pea10d]. Variables [Ber18, Iss18a, Iss18b]. Variate [CP14, RS15]. Variates [PY18]. Variation [CK15, Gri19, Pea12c, Sim14]. Variations [Emm13]. Various [PY18]. Vespa [Lat14, TBP11]. View [Las12]. Viscera [GB13]. Volume [Ano11c, Ano13f, Ano14b, Ano15c, Ano18]. Vulgaris [Lat14, TBP11].

W. [Ano15a, Ker17]. Webb [Pea12a]. Weight [Eld12, Eld14a, GB13, Har13b, Har14b]. Weldon [Ano15a]. Which [Pea15a, Stu17, TM12]. Who [Pea14c]. Whorl [Har11]. Within [Har14a]. Without [YP16]. Women [Eld14b]. Work [Pea14d]. Workers [Lat14, Sno12b].

Young [Jen12].

Zaaijer [Ano13c, Smi12].

References

Acton:1912:INE

[AH12] Hugh W. Acton and W. F. Harvey. The increase in the number of erythrocytes with altitude. Biometrika, 8(3/4):280–291, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331583.

Adler:1910:HEC

[AM10] J. E. Adler and J. McIntosh. Histological examination of a case of albinism. Biometrika, 7(3):237–243, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345383.

Anderson:1914:NES

[And14] O. Anderson. Nochmals über “The Elimination of Spurious Correlation Due to Position in Time or Space (German) [Comments on …]. Biometrika, 10(2/3):269–279, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331784.

Anonymous:1910:B

[Ano10] Anonymous. Bibliography. Biometrika, 7(3):414–420, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345399.

Anonymous:1911:FG

[Ano11a] Anonymous. Francis Galton. Biometrika, 8(1/2):??, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331435.

Anonymous:1911:FM

[Ano11b] Anonymous. Front matter. Biometrika, 8(1/2):ii, July 1911. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331434.

Anonymous:1911:VI

[Ano11c] Anonymous. Volume information. Biometrika, 8(1/2):i–vii, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331433.

Anonymous:1913:B

[Ano13a] Anonymous. Bibliography. Biometrika, 9(1/2):140–158, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331804.

Anonymous:1913:CAM

[Ano13b] Anonymous. Corrigenda: “Anthropometry of Modern Egyptians”. Biometrika, 9(1/2):332, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331811. See [Cra11].

Anonymous:1913:COO

[Ano13c] Anonymous. Corrigenda: “Observations on the Occipital Bone in a Series of Egyptian Skulls, with Especial Reference to the Persistence of the Synchondrosis Condylo-Squamosa (Zaaijer; Synchondrosis Intraoccipitalis Posterior, B.N.A.)”. Biometrika, 9(1/2):332, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331810. See

[Smi12].

Anonymous:1913:NRR

[Ano13d] Anonymous. Note on a ready reckoner for the obstetrician. Biometrika, 9(3/4):539–540, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331908.

Anonymous:1913:PEF

[Ano13e] Anonymous. On the probable errors of frequency constants. Biometrika, 9(1/2):1–10, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331796.

Anonymous:1913:VI

[Ano13f] Anonymous. Volume information. Biometrika, 9(1/2):i–vii, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331795.

Anonymous:1914:A

[Ano14a] Anonymous. Announcement. Biometrika, 10(1):196, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331751.

Anonymous:1914:VI

[Ano14b] Anonymous. Volume information. Biometrika, 10(1):i–vi, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331736.

Anonymous:1915:ABW

[Ano15a] Anonymous. Appendix to Biometrika: W. F. R. Weldon’s mice breeding experiments. Records of matings. Biometrika, 11(1/2): 145+1–60, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331887.

Anonymous:1915:EDS

[Ano15b] Anonymous. Editorial: On the distribution of the standard deviations of small samples: Appendix I. To papers by “Student” and R. A. Fisher. Biometrika, 10(4):522–529, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331839.

Anonymous:1915:VI

[Ano15c] Anonymous. Volume information. Biometrika, 11(1/2):i–vi, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331879.

Anonymous:1917:PEM

[Ano17] Anonymous. The probable error of a Mendelian class frequency. Biometrika, 11(4):429–432, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331834.

Anonymous:1918:VI

[Ano18] Anonymous. Volume information. Biometrika, 12(1/2):i–iv, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331928.

Anonymous:1919:P

[Ano19] Anonymous. Peccavimus. Biometrika, 12(3/4):259–281, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331770.

Bell:1911:PBR

[Bel11] Julia Bell. On pulse and breathing rates and their relation to stature. Biometrika, 8(1/2):232–236, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331447.

Bergstrom:1918:MFC

[Ber18] Sverker Bergstrom. Sur les moments de la fonction de correlation normale de n variables. (French) [On the moments of the normal correlation function of n variables]. Biometrika, 12(1/2):177–183, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331935.

Benington:1911:CTC

[BP11] R. Crewdson Benington and Karl Pearson. Cranial type-contours. Biometrika, 8(1/2):123–201, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331444.

Benington:1912:SNS

[BP12] R. Crewdson Benington and Karl Pearson. A study of the Negro skull with special reference to the Congo and gaboon crania. Biometrika, 8(3/4):292–339, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331584.

Brown:1910:OSM

[Bro10] William Brown. An objective study of mathematical intelligence. Biometrika, 7(3):352–367, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345389.

Cool:1915:VCN

[CK15] C. Cool and A. N. Koopmans. Variation and correlation of the number of umbel rays of some Umbelliferae. Biometrika, 11(1/2):38–49, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331881.

Claremont:1916:CBC

[Cla16] C. A. Claremont. On the correlation between the “corrected” cancer and diabetes deathrates. Biometrika, 11(3):191–200, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331847.

Cockayne:1914:PF

[Coc14] E. A. Cockayne. A piebald family. Biometrika, 10(2/3):197–200, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331780.

Cave:1914:NIV

[CP14] Beatrice M. Cave and Karl Pearson. Numerical illustrations of the variate difference correlation method. Biometrika, 10(2/3):340–355, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331787.

Craig:1911:AME

[Cra11] J. I. Craig. Anthropometry of modern Egyptians. Biometrika, 8(1/2):66–78, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2331440. See corrigenda [Ano13b].

Davenport:1910:DGC

[Dav10] C. B. Davenport. Dr. Galloway’s “canary breeding”. Biometrika, 7(3):398–400, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345392.

DeSouza:1913:MPS

[De 13] D. H. De Souza. The measurements of the pelvis, with special reference to obstetric prediction. Biometrika, 9(3/4):486–529, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331903.

Doodson:1917:RMM

[Doo17] Arthur T. Doodson. Relation of the mode, median and mean in frequency curves. Biometrika, 11(4):425–429, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331833.

Duncker:1911:NHV

[Dun11] Georg Duncker. Note on high values of β_{1} and β_{2}. Biometrika, 8(1/2):238, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331449.

E:1910:BRBa

[E.10a] W. P. E. Book review: Interpolationsrechnung by T. N. Thiele. Biometrika, 7(3):413, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345397.

E:1910:BRBb

[E.10b] W. P. E. Book review: The Theory of the Construction of Tables of Mortality by G. F. Hardy. Biometrika, 7(3):413–414, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345398.

E:1912:AAP

[E.12] W. P. E. An attempt to ascertain the prevalence of syphilis in a large urban population. Biometrika, 8(3/4):437, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331595.

E:1913:BRB

[E.13] W. P. E. Book review: Studies in the Meaning and Relationships of Birth and Death Rates. I. The Relationship Between “Corrected” Death-Rates and Life Table Death-Rates by John Brownlee. Biometrika, 9(3/4):537, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331906.

Elderton:1912:RSW

[Eld12] Ethel M. Elderton. On the relation of stature and weight to pigmentation. Biometrika, 8(3/4):340–353, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331585.

Elderton:1914:HWS

[Eld14a] Ethel M. Elderton. Height and weight of school children in Glasgow. Biometrika, 10(2/3):288–339, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331786.

Elderton:1914:NIM

[Eld14b] Ethel M. Elderton. Note on infantile mortality and employment of women. Biometrika, 10(1):193–196, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331750.

Elderton:1919:LHA

[Eld19] Ethel M. Elderton. Life-history albums. Biometrika, 12(3/4): 373–374, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331778.

Emmons:1913:SVF

[Emm13] Arthur Brewster Emmons. A study of the variations in the female pelvis, based on observations made on 217 specimens of the American Indian squaw. Biometrika, 9(1/2):34–57, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331799.

Elderton:1915:FEN

[EP15a] Ethel M. Elderton and Karl Pearson. Further evidence of natural selection in man. Biometrika, 10(4):488–506, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331837.

Elderton:1915:IID

[EP15b] Ethel M. Elderton and Karl Pearson. The influence of isolation on the diphtheria attack- and death-rates. Biometrika, 10(4):549–569, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331841.

Everitt:1910:TTF

[Eve10] P. F. Everitt. Tables of the tetrachoric functions for fourfold correlation tables. Biometrika, 7(4):437–451, November 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2345377.

Everitt:1912:STF

[Eve12] P. F. Everitt. Supplementary tables for finding the correlation coefficient from tetrachoric groupings. Biometrika, 8(3/4):385–395, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331587.

Everitt:1919:QC

[Eve19] P. F. Everitt. Quadrature coefficients. Biometrika, 12(3/4):282–283, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331771.

Fisher:1915:FDV

[Fis15] R. A. Fisher. Frequency distribution of the values of the correlation coefficient in samples from an indefinitely large population. Biometrika, 10(4):507–521, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331838.

Galloway:1910:CBR

[Gal10] A. Rudolf Galloway. Canary breeding. A rejoinder to C. B. Davenport. Biometrika, 7(3):401–403, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345393.

Galloway:1912:HC

[Gal12a] A. Rudolf Galloway. Hybridisation of canaries. Biometrika, 8 (3/4):435–436, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331594.

Galloway:1912:NPH

[Gal12b] A. Rudolf Galloway. Notes on the pigmentation of the human Iris. Biometrika, 8(3/4):267–279, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331582.

Greenwood:1913:SSW

[GB13] M. Greenwood, Jr. and J. W. Brown. A second study of the weight, variability and correlation of the human viscera. Biometrika, 9(3/4):473–485, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331902.

Goring:1919:CBG

[Gor19] Katie MacDonald Goring. Charles B. Goring 1870–1919. Biometrika, 12(3/4):297–307, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331773.

Gilby:1911:STA

[GP11] Walter H. Gilby and Karl Pearson. On the significance of the Teacher’s appreciation of general intelligence. Biometrika, 8(1/2): 94–108, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331442.

Greenwood:1913:ERS

[Gre13] M. Greenwood, Jr. On errors of random sampling in certain cases not suitable for the application of a “normal” curve of frequency. Biometrika, 9(1/2):69–90, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331801.

Grier:1919:VDL

[Gri19] N. M. Grier. Variation and distribution of leaves in sassafras. Biometrika, 12(3/4):372–373, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331777.

Greenwood:1910:BSP

[GW10] M. Greenwood, Jr. and J. D. C. White. A biometric study of phagocytosis with special reference to the “opsonic index”. Second memoir. On the distribution of the means of samples. Biometrika, 7(4):505–530, November 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345379.

Harris:1910:SEO

[Har10a] J. Arthur Harris. On the selective elimination occurring during the development of the fruits of Staphylea. Biometrika, 7(4):452–504, November 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2345378.

Harris:1910:QSM

[Har10b] J. Arthur Harris. A quantitative study of the morphology of the fruit of the bloodroot, Sanguinaria canadensis. Biometrika, 7(3):305–351, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345388.

Harris:1911:CBS

[Har11] J. Arthur Harris. On the correlation between somatic characters and fertility: Illustrations from the involucral whorl of hibiscus. Biometrika, 8(1/2):52–65, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331439.

Harris:1913:CIC

[Har13a] J. Arthur Harris. On the calculation of intra-class and inter-class coefficients of correlation from class moments when the number of possible combinations is large. Biometrika, 9(3/4):446–472, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331901.

Harris:1913:RBW

[Har13b] J. Arthur Harris. The relationship between the weight of the seed planted and the characteristics of the plant produced. I. Biometrika, 9(1/2):11–21, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331797.

Harris:1914:SVI

[Har14a] J. Arthur Harris. On spurious values of intra-class correlation coefficients arising from disorderly differentiation within the classes. Biometrika, 10(2/3):412–416, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331791.

Harris:1914:RBW

[Har14b] J. Arthur Harris. The relationship between the weight of the seed planted and the characteristics of the plant produced. II. Biometrika, 10(1):72–84, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331740.

Harris:1916:CPH

[Har16] J. Arthur Harris. A contribution to the problem of homotyposis: Data from the legume cercis canadensis. Biometrika, 11(3):201–214, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331848.

Heron:1910:ADP

[Her10a] David Heron. An abac for determining the probable errors of correlation coefficients. Biometrika, 7(3):411, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2345395.

Heron:1910:ICS

[Her10b] David Heron. Inheritance in canaries: A study in Mendelism. Biometrika, 7(3):403–410, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345394.

Heron:1910:PEP

[Her10c] David Heron. On the probable error of a partial correlation coefficient. Biometrika, 7(3):411–412, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345396.

Heron:1911:DCF

[Her11] David Heron. The danger of certain formulae suggested as substitutes for the correlation coefficient. Biometrika, 8(1/2):109–122, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331443.

Heron:1914:ESR

[Her14a] David Heron. An examination of some recent studies of the inheritance factor in insanity. Biometrika, 10(2/3):356–383, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331788.

Heron:1914:NRS

[Her14b] David Heron. Note on reproductive selection. Biometrika, 10 (2/3):419–420, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331794.

Hirshberg:1913:SIB

[Hir13] Leonard Keene Hirshberg. Selection and intermediates in Bacillus Coli. Biometrika, 9(1/2):331–332, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331809.

Hosny:1914:SNI

[Hos14] M. Hosny. Statistical notes on the influence of education in Egypt. Biometrika, 10(2/3):280–287, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331785.

Isserlis:1914:FDC

[Iss14a] L. Isserlis. Formulae for the determination of the capacity of the Negro skull from external measurements. Biometrika, 10(1):188–193, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331748.

Isserlis:1914:PCR

[Iss14b] L. Isserlis. On the partial correlation ratio. Biometrika, 10 (2/3):391–411, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331790.

Isserlis:1915:PCR

[Iss15] L. Isserlis. On the partial correlation-ratio: Part II. Numerical. Biometrika, 11(1/2):50–66, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331882.

Isserlis:1916:CPE

[Iss16] L. Isserlis. On certain probable errors and correlation coefficients of multiple frequency distributions with skew regression. Biometrika, 11(3):185–190, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331846.

Isserlis:1917:RSD

[Iss17] L. Isserlis. On the representation of statistical data. Biometrika, 11(4):418–425, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331832.

Isserlis:1918:FDM

[Iss18a] L. Isserlis. Formulae for determining the mean values of products of deviations of mixed moment coefficients in two to eight variables in samples taken from a limited population. Biometrika, 12 (1/2):183–184, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331936.

Isserlis:1918:FPM

[Iss18b] L. Isserlis. On a formula for the product-moment coefficient of any order of a normal frequency distribution in any number of variables. Biometrika, 12(1/2):134–139, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331932.

Jenkinson:1912:GVC

[Jen12] J. W. Jenkinson. Growth, variability and correlation in Young trout. Biometrika, 8(3/4):444–455, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331598.

Jones:1910:VTO

[Jon10] H. Gertrude Jones. On the value of the teachers’ opinion of the general intelligence of school children. Biometrika, 7(4):542–548, November 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2345381.

Ker:1917:WRM

[Ker17] W. P. Ker. W. R. Macdonell. Biometrika, 11(4):281–283, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331826.

Laski:1912:MVR

[Las12] H. J. Laski. A Mendelian view of racial heredity. Biometrika, 8(3/4):424–430, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331592.

Latter:1914:CMQ

[Lat14] Oswald H. Latter. Clypeal markings of queens, drones and workers of Vespa Vulgaris. Biometrika, 10(2/3):201–207, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331781.

Lee:1914:TGT

[Lee14] Alice Lee. Table of the Gaussian “tail” functions; when the “tail” is larger than the body. Biometrika, 10(2/3):208–214, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331782.

Lee:1915:TS

[Lee15] Alice Lee. Tuberculosis and segregation. Biometrika, 10(4):530–548, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331840.

Lee:1917:FST

[Lee17] Alice Lee. Further supplementary tables for determining high correlations from tetrachoric groupings. Biometrika, 11(4):284–291, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331827.

Lloyd:1911:IF

[Llo11] R. E. Lloyd. The inheritance of fertility. Biometrika, 8(1/2):244–247, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331451. See

remarks [Pea11e].

Macdonald:1911:PHE

[Mac11] David Macdonald. Pigmentation of the hair and eyes of children suffering from the acute fevers, its effect on susceptibility, recuperative power and race selection. Biometrika, 8(1/2):13–39, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331437.

Macdonell:1913:ELA

[Mac13] W. R. Macdonell. On the expectation of life in Ancient Rome, and in the provinces of Hispania and Lusitania, and Africa. Biometrika, 9(3/4):366–380, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331898.

Maynard:1910:SSC

[May10] G. D. Maynard. A statistical study in cancer death-rates. Biometrika, 7(3):276–304, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345387.

Maynard:1911:NAD

[May11] G. D. Maynard. A note on the age distribution of deaths from diabetes mellitus. Biometrika, 8(1/2):225–231, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331446.

Maynard:1914:NNP

[May14] C. D. Maynard. Note on a Negro piebald. Biometrika, 10(1):193, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331749.

McKendrick:1912:SMC

[McK12] A. G. McKendrick. A simplified method of calculating frequencies of occurrence, from a large number of unequal probabilities. Biometrika, 8(3/4):413–419, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331590.

McMullan:1913:IDKb

[MP13a] George McMullan and Karl Pearson. On the inheritance of deformity known as split-foot or lobster-claw: Addendum. Biometrika, 9(3/4):540, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331909.

McMullan:1913:IDKa

[MP13b] George McMullan and Karl Pearson. On the inheritance of the deformity known as split-foot or lobster-claw. Biometrika, 9 (3/4):381–390, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331899.

Orensteen:1915:CAM

[Ore15] Myer M. Orensteen. Correlation of anthropometrical measurements in Cairo-born natives. Biometrika, 11(1/2):67–81, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331883.

P:1912:FPS

[P.12] K. P. On a fallacious proof of Sheppard’s correction. Biometrika, 8(3/4):443, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331597.

P:1914:CMM

[P.14] K. P. Correction of a misstatement by Mr Major Greenwood, Junior. Biometrika, 10(2/3):418–419, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331793.

P:1915:MME

[P.15] K. P. Measurements of medieval English femora. Biometrika, 10(4):573–574, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331843.

Paton:1914:SSD

[Pat14] D. Noël Paton. The statistical study of dietaries, a reply to Professor Karl Pearson. Biometrika, 10(1):169–172, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331743. See [Pea13e].

Pearson:1913:HCG

[PE13] Karl Pearson and Ethel M. Elderton. On the hereditary character of general health. Biometrika, 9(1/2):320–329, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331807.

Pearson:1910:DBS

[Pea10a] Karl Pearson. Darwinism, biometry and some recent biology. I. Biometrika, 7(3):368–385, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345390.

Pearson:1910:NIA

[Pea10b] Karl Pearson. Note on internal albinism. Biometrika, 7(3):244–247, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345384.

Pearson:1910:NSI

[Pea10c] Karl Pearson. Note on the separate inheritance of quantity and quality in cows’ milk. Biometrika, 7(4):548–550, November 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345382.

Pearson:1910:NMD

[Pea10d] Karl Pearson. On a new method of determining correlation, when one variable is given by alternative and the other by multiple categories. Biometrika, 7(3):248–257, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345385.

Pearson:1910:CID

[Pea10e] Karl Pearson. On the constants of index-distributions as deduced from the like constants for the components of the ratio, with special reference to the opsonic index. Biometrika, 7(4):531–541, November 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345380.

Pearson:1910:EDF

[Pea10f] Karl Pearson. On the effect of differential fertility of degeneracy. Biometrika, 7(3):258–275, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345386.

Pearson:1911:FRL

[Pea11a] Karl Pearson. Further remarks on the law of ancestral heredity. Biometrika, 8(1/2):239–243, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331450.

Pearson:1911:CMC

[Pea11b] Karl Pearson. On a correction to be made to the correlation ratio η. Biometrika, 8(1/2):254–256, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331454.

Pearson:1911:PTI

[Pea11c] Karl Pearson. On the probability that two independent distributions of frequency are really samples from the same population. Biometrika, 8(1/2):250–254, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331453.

Pearson:1911:OIM

[Pea11d] Karl Pearson. The opsonic index — “mathematical error and functional error”. Biometrika, 8(1/2):203–224, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331445.

Pearson:1911:RPL

[Pea11e] Karl Pearson. Remarks on Professor Lloyd’s note on inheritance of fertility. Biometrika, 8(1/2):247–249, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331452. See [Llo11].

Pearson:1912:CHD

[Pea12a] Karl Pearson. On “Cancer Houses,” from the data of the late Th. Law Webb, M.D. Biometrika, 8(3/4):430–435, January 1912. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331593.

Pearson:1912:AMC

[Pea12b] Karl Pearson. On the appearance of multiple cases of disease in the same house. Biometrika, 8(3/4):404–412, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331589.

Pearson:1912:GTI

[Pea12c] Karl Pearson. On the general theory of the influence of selection on correlation and variation. Biometrika, 8(3/4):437–443, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331596.

Pearson:1913:NHP

[Pea13a] Karl Pearson. Note on the Honduras piebald. Biometrika, 9 (1/2):330–331, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331808.

Pearson:1913:NSC

[Pea13b] Karl Pearson. Note on the surface of constant association. Biometrika, 9(3/4):534–537, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331905.

Pearson:1913:MIB

[Pea13c] Karl Pearson. On the measurement of the influence of “broad categories” on correlation. Biometrika, 9(1/2):116–139, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331803.

Pearson:1913:PEC

[Pea13d] Karl Pearson. On the probable error of a coefficient of correlation as found from a fourfold table. Biometrika, 9(1/2):22–33, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331798.

Pearson:1913:SSD

[Pea13e] Karl Pearson. The statistical study of dietaries. Biometrika, 9 (3/4):530–533, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2331904. See reply [Pat14].

Pearson:1914:NEC

[Pea14a] K. Pearson. Note on the essential conditions that a population breeding at random should be in a stable state. Biometrika, 10(1):175–178, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331745.

Pearson:1914:EMC

[Pea14b] Karl Pearson. On an extension of the method of correlation by grades or ranks. Biometrika, 10(2/3):416–418, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331792.

Pearson:1914:CER

[Pea14c] Karl Pearson. On certain errors with regard to multiple correlation occasionally made by those who have not adequately studied this subject. Biometrika, 10(1):181–187, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331747.

Pearson:1914:PTI

[Pea14d] Karl Pearson. On the probability that two independent distributions of frequency are really samples of the same population, with special reference to recent work on the identity of trypanosome strains. Biometrika, 10(1):85–143, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331741.

Pearson:1914:SSD

[Pea14e] Karl Pearson. The statistical study of dietaries. A rejoinder. Biometrika, 10(1):172–174, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331744.

Pearson:1915:CTC

[Pea15a] Karl Pearson. On certain types of compound frequency distributions in which the components can be individually described by binomial series. Biometrika, 11(1/2):139–144, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331886.

Pearson:1915:PEC

[Pea15b] Karl Pearson. On the probable error of a coefficient of mean square contingency. Biometrika, 10(4):570–573, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331842.

Pearson:1915:PSO

[Pea15c] Karl Pearson. On the problem of sexing osteometric material. Biometrika, 10(4):479–487, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331836.

Pearson:1916:SNP

[Pea16a] Karl Pearson. On some novel properties of partial and multiple correlation coefficients in a universe of manifold characteristics. Biometrika, 11(3):231–238, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331850.

Pearson:1916:AGF

[Pea16b] Karl Pearson. On the application of “goodness of fit” tables to test regression curves and theoretical curves used to describe observational or experimental data. Biometrika, 11(3):239–261, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331851.

Pearson:1916:GTM

[Pea16c] Karl Pearson. On the general theory of multiple contingency with special reference to partial contingency. Biometrika, 11(3):145–158, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331844.

Pearson:1917:PEB

[Pea17] Karl Pearson. On the probable error of biserial η. Biometrika, 11(4):292–302, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331828.

Pearson:1919:CFB

[Pea19a] Karl Pearson. The check to the fall in the Phthisis death-rate since the discovery of the tubercle bacillus and the adoption of modern treatment. Biometrika, 12(3/4):374–376, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331779.

Pearson:1919:IPC

[Pea19b] Karl Pearson. Inheritance of psychical characters. Biometrika, 12 (3/4):367–372, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331776.

Pearson:1919:GTT

[Pea19c] Karl Pearson. On generalised Tchebycheff theorems in the mathematical theory of statistics. Biometrika, 12(3/4):284–296, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331772.

Pearson:1913:TA

[PH13] Karl Pearson and David Heron. On theories of association. Biometrika, 9(1/2):159–315, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331805.

Pairman:1919:CMC

[PP19] Eleanor Pairman and Karl Pearson. On corrections for the momentcoefficients of limited range frequency distributions when there are finite or infinite ordinates and any slopes at the terminals of the range. Biometrika, 12(3/4):231–258, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331769.

Pearson:1916:CED

[PT16] Karl Pearson and J. F. Tocher. On criteria for the existence of differential deathrates. Biometrika, 11(3):159–184, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331845.

Pearson:1918:PMV

[PY18] K. Pearson and A. W. Young. On the product-moments of various orders of the normal correlation surface of two variates. Biometrika, 12(1/2):86–92, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331930.

Ryley:1913:SNB

[RBP13] Kathleen V. Ryley, Julia Bell, and Karl Pearson. A study of the nasal bridge in the anthropoid apes and its relationship to the nasal bridge in man. Biometrika, 9(3/4):391–445, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331900.

Rhind:1910:ATD

[Rhi10] A. Rhind. Additional tables and diagram for the determination of the errors of type of frequency distributions. Second paper. Biometrika, 7(3):386–397, April 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345391.

Rock:1911:TBD

[Roc11] Frank Rock. The tooth-brush and decayed teeth. Biometrika, 8 (1/2):237–238, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331448.

Rowan:1914:HAC

[RPB14] William Rowan, K. M. Parker, and Julia Bell. On homotyposis and allied characters in eggs of the common tern. Biometrika, 10(1):144– 168, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331742.

Ritchie-Scott:1915:NPE

[RS15] A. Ritchie-Scott. Note on the probable error of the coefficient of correlation in the variate difference correlation method. Biometrika, 11(1/2):136–138, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331885.

Ritchie-Scott:1918:CCP

[RS18] A. Ritchie-Scott. The correlation coefficient of a polychoric table. Biometrika, 12(1/2):93–133, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331931.

Rowan:1919:NEC

[RWS^{+}19] W. Rowan, E. Wolff, P. L. Sulman, K. Pearson, E. Isaacs, E. M. Elderton, and M. Tildesley. On the nest and eggs of the common tern (S. Fluviatilis). A cooperative study. Biometrika, 12(3/4):308–354, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331774.

Saunders:1912:PRS

[Sau12] A. M. Carr Saunders. Pigmentation in relation to selection and to anthropometric characters. Biometrika, 8(3/4):354–384, January 1912.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331586.

Schuster:1911:FRO

[Sch11] E. Schuster. First results from the Oxford anthropometric laboratory. Biometrika, 8(1/2):40–51, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331438.

Simpson:1914:CSS

[Sim14] James J. Simpson. Contribution to a statistical study of the Cruciferæ. Variation in the flowers of Lepidium Draba Linnæus. Biometrika, 10(2/3):215–268, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331783.

Smith:1912:OOB

[Smi12] H. Dorothy Smith. Observations on the occipital bone in a series of Egyptian skulls, with especial reference to the persistence of the Synchondrosis Condylo-Squamosa (Zaaijer; Synchondrosis Intraoccipitalis Posterior, B.N.A.). Biometrika, 8(3/4):257–266, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331581. See corri-

genda [Ano13c].

Smith:1916:BVC

[Smi16] Kirstine Smith. On the ‘best’ values of the constants in frequency distributions. Biometrika, 11(3):262–276, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331852.

Smith:1918:SDA

[Smi18] Kirstine Smith. On the standard deviations of adjusted and interpolated values of an observed polynomial function and its constants and the guidance they give towards a proper choice of the distribution of observations. Biometrika, 12(1/2):1–85, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331929.

Snow:1912:ACC

[Sno12a] E. C. Snow. The application of the correlation coefficient to Mendelian distributions. Biometrika, 8(3/4):420–424, January 1912.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331591.

Snow:1912:BWS

[Sno12b] E. C. Snow. Biometric workers and statistical reviewers. Biometrika, 8(3/4):456–460, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331599.

Snow:1913:INS

[Sno13] E. C. Snow. The intensity of natural selection in man. Biometrika, 9(1/2):58–68, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331800.

Soper:1913:PEC

[Sop13] H. E. Soper. On the probable error of the correlation coefficient to a second approximation. Biometrika, 9(1/2):91–115, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331802.

Soper:1914:PEB

[Sop14a] H. E. Soper. On the probable error of the bi-serial expression for the correlation coefficient. Biometrika, 10(2/3):384–390, November 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331789.

Soper:1914:TPE

[Sop14b] H. E. Soper. Tables of Poisson’s exponential binomial limit. Biometrika, 10(1):25–35, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331738.

Stannus:1913:APA

[Sta13] Hugh Stannus Stannus. Anomalies of pigmentation among natives of Nyasaland: A contribution to the study of albinism. Biometrika, 9(3/4):333–365, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331897.

Stannus:1914:CAN

[Sta14] Hugh Stannus Stannus. Congenital anomalies in a native African race. Biometrika, 10(1):1–24, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331737.

Student:1913:CMC

[Stu13] Student. The correction to be made to the correlation ratio for grouping. Biometrika, 9(1/2):316–320, March 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331806.

Student:1914:ESC

[Stu14] Student. The elimination of spurious correlation due to position in time or space. Biometrika, 10(1):179–180, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331746.

Student:1917:TEP

[Stu17] Student. Tables for estimating the probability that the mean of a unique sample of observations lies between −∞ and any given distance of the mean of the population from which the sample is drawn. Biometrika, 11(4):414–417, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331831.

Student:1919:EDP

[Stu19] Student. An explanation of deviations from Poisson’s law in practice. Biometrika, 12(3/4):211–215, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331767.

Soper:1917:DCC

[SYC^{+}17] H. E. Soper, A. W. Young, B. M. Cave, A. Lee, and K. Pearson. On the distribution of the correlation coefficient in small samples. Appendix II to the papers of “Student” and R. A. Fisher. Biometrika, 11(4):328–413, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331830.

Thomson:1911:TCS

[TBP11] E. Y. Thomson, Julia Bell, and Karl Pearson. A third cooperative study of Vespa Vulgaris. Comparison of queens of a single nest with queens of the general autumn population. Biometrika, 8(1/2):1–12, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331436.

Tchouproff:1918:MEM

[Tch18] Al. A. Tchouproff. On the mathematical expectation of the moments of frequency distributions. Biometrika, 12(1/2):140–169, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331933.

Tchouproff:1919:MEM

[Tch19] Al. A. Tchouproff. On the mathematical expectation of the moments of frequency distributions. Biometrika, 12(3/4):185–210, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331766.

Thomson:1915:SCM

[Tho15] Eveline Y. Thomson. A study of the crania of the Moriori, or Aborigines of the Chatham Islands, now in the museum of the Royal College of Surgeons. Biometrika, 11(1/2):82–135, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331884.

Thomson:1919:CGF

[Tho19a] Godfrey H. Thomson. The criterion of goodness of fit of psychophysical curves. Biometrika, 12(3/4):216–230, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331768.

Thomson:1919:DPH

[Tho19b] Godfrey H. Thomson. On the degree of perfection of hierarchical order among correlation coefficients. Biometrika, 12(3/4):355–366, November 1919. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331775.

Tildesley:1918:PNA

[Til18] M. L. Tildesley. Preliminary note on the association of steadiness and rapidity of hand with artistic capacity. Biometrika, 12(1/2):170–177, November 1918. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331934.

Troup:1912:NEW

[TM12] J. McD. Troup and G. D. Maynard. Note on the extent to which the distribution of cases of disease in houses is determined by the laws of chance. Biometrika, 8(3/4):396–403, January 1912. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331588.

Turnbull:1913:CFM

[Tur13] Hubert M. Turnbull. A case of familial malformation in a Fowl’s head. Biometrika, 9(3/4):538–539, October 1913. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331907.

Waite:1910:MMS

[Wai10] H. Waite. Mosquitoes and malaria. A study of the relation between the number of mosquitoes in a locality and the malaria rate. Biometrika, 7(4):421–436, November 1910. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345376.

Waite:1911:TEG

[Wai11] H. Waite. The Teacher’s estimation of the general intelligence of school children. Biometrika, 8(1/2):79–93, July 1911. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331441.

Waite:1915:AFP

[Wai15] H. Waite. Association of finger-prints. Biometrika, 10(4):421–478, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331835.

Warren:1917:PRS

[War17] Ernest Warren. A preliminary report on some breeding experiments with foxgloves. Biometrika, 11(4):303–327, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331829.

Whitaker:1914:PLS

[Whi14] Lucy Whitaker. On the Poisson law of small numbers. Biometrika, 10(1):36–71, April 1914. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331739.

Whiting:1915:ATP

[Whi15] Madeline H. Whiting. On the association of temperature, pulse and respiration with physique and intelligence in criminals: A study in criminal anthropometry. Biometrika, 11(1/2):1–37, November 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331880.

Young:1916:NSD

[You16] Andrew W. Young. Note on the standard deviations of samples of two or three. Biometrika, 11(3):277–280, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331853.

Young:1916:PEC

[YP16] Andrew W. Young and Karl Pearson. On the probable error of a coefficient of contingency without approximation. Biometrika, 11(3):215–230, May 1916. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331849.